# The DREAMS Project: Disentangling the Impact of Halo-to-Halo Variance and Baryonic Feedback on Milky Way Dark Matter Density Profiles

### v1

@@ 4.2.1 Comparison with Bursty Feedback Model
It is not clear, however, whether bursty feedback disrupting the potential and preventing adiabatic contraction is a generic feature of FIRE-2 or a feature of the specific implementation of the model. An analogous suite of Milky Way-mass halos within the FIRE-2 model (or its successor FIRE-3; Hopkins et al. 2023) would be particularly valuable for constraining this. Such a comparison would help determine whether the deviations from adiabatic contraction identified by Hussein et al. (2025) are a generic outcome of bursty feedback or depend sensitively on the details of how feedback is modeled in FIRE.
5 Conclusion
@@
  • Overall, there are only minor variations in the density profiles of Milky Way-mass halos with variations to the IllustrisTNG model (Section 3.1 Figure 2). Virtually all variation in the density profiles is driven by halo-to-halo variation, with a relatively small contribution from the two supernova parameters that only create differences of $\pm\sim 20-50\%$ in the inner regions ($\lesssim 0.01R_{200}$) of the halos.
@@
  • We fit the density profiles with a generalized NFW profile (e.g., Jaffe, 1983; Hernquist, 1990) which has two normalization parameters and three shape parameters (Equation 9, Section 3.2). Again, intrinsic halo-to-halo variation plays the dominant role in setting both the normalization and shape of halos. Beyond the halo-to-halo variance, changes to the supernova wind speed ($\kappa_{w}$) can also vary the normalization of the profile (Figure 3) and increases to the wind energies ($\bar{e}_{w}$) can drive the shape of the profiles closer to NFW (Figure 4). Physically, this suggests that ejecting faster winds more strongly disrupts the potential, whereas more wind energy effectively suppresses stellar mass growth.
@@
  • We quantified how much each halo mass increases using the central mass growth: the ratio of the DMO simulation enclosed mass to the hydro simulation enclosed mass at $0.01R_{200}$ (Figure 5). Here, the astrophysics variations play a dominant role. Larger values of $\bar{e}_{w}$ and $\kappa_{w}$ are more efficient at preventing stellar mass growth (left two panels of Figure 6), thus the halo contracts less. On the other hand, we find the opposite trend in $\epsilon_{f,\,\mathrm{high}}$: increasing black hole feedback increases the work required to contract (right panel of Figure 6, see also Appendix C).
@@
  • We compared the results of the hydro simulations to DMO simulations with matched initial conditions (Section 4.1). On the whole, the DMO simulations have a lower density in the inner regions ($\lesssim 0.01R_{200}$) than the hydro simulations (Figure 7). The extent to which the density is enhanced depends on the feedback, however. In fact, higher values of $\bar{e}_{w}$ tend towards a ratio of unity with respect to the DMO simulations, in good agreement with the results of the profile shape and central mass growth.
@@
  • We quantified whether or not our halos are consistent with approximations to adiabatic contraction (Section 4.2). Halos generally agree within $\sim 10\%$ regardless of the feedback or cosmology (Figure 8); however, systems with late-time mergers that occur at slightly different times between the hydro and DMO simulations tend to disagree significantly. As with many previous results, the supernova wind energy changes the picture due to the lack of baryon mass accumulation: these halos contract adiabatically.Finally, we showed that the bursty feedback FIRE(-2) model produces significantly larger deviations from adiabatic contraction than the average DREAMS halos (Figure 9). This suggests that what drives deviations from adiabatic contractions are bursts of stellar feedback, not just strong feedback.
@@
Taken together, these results show that halo-to-halo variance is the dominant source of scatter in Milky Way–mass dark matter profiles in the TNG model, with $\bar{e}_{w}$ and $\kappa_{w}$ feedback parameters producing secondary but non-negligible effects. AGN feedback plays a smaller role, and cosmological parameter variations are essentially irrelevant at this mass scale.
@@
Raw data from the DREAMS CDM and WDM suites are available upon reasonable request of the authors. The scripts used to generate the reduced data and all other scripts used to support the findings of this work are made publicly available here: github.com/AlexGarcia623/DREAMS-Density-Profiles
@@ Acknowledgements
AMG, PT, JK, NA, AB, XO, AF, and NK acknowledge support from the National Science Foundation under Cooperative Agreement 2421782 and the Simons Foundation grant MPS-AI-00010515 awarded to NSF-Simons AI Institute for Cosmic Origins – CosmicAI, https://www.cosmicai.org/. ML acknowledges support from the Simons Investigator Award. AC acknowledges the hospitality of the Weizmann Institute of Science and the support from the Benoziyo Endowment Fund for the Advancement of Science. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (grant agreement No. 101117510). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union. The European Union cannot be held responsible for them. LN is supported by the Sloan Fellowship, the NSF CAREER award 2337864, NSF award 2307788, and by the NSF award PHY2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) XS acknowledges the support of the NASA theory grant JWST-AR-04814. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. The authors also thank the Simons Foundation for their support in hosting and organizing workshops on the DREAMS Project.
@@ Appendix A Emulator Validation Test
As mentioned in Section 2.3, we build a series of neural network emulators to help support the findings of this work. We validate each emulator (Figures 3456C1, and D1) by comparing its predictions to the raw simulation outputs (following Rose et al., 2025a). We show Figure A1 as an example of this validation process. The left-hand panel of Figure A1 is identical to Figure 4 with the addition of the true rolling mean (solid black line) and standard deviation (gray shaded region) from the measured simulation data points. The raw simulation results largely track the emulator ones, although there is a small shift in the outer slope. One potential reason for this is that the emulator is trained to predict single parameter dependencies (i.e., we keep all other parameters fixed), whereas the raw simulation outputs include inherent co-variation of parameters. To capture this co-variance, we would need a different architecture, such as normalizing flows (see, e.g., E. Lilie In Preparation). Regardless, the emulated prediction for the mean is largely consistent with these raw simulation outputs even with our simple emulator. Figure A1 further demonstrates the utility of the emulator: the true rolling mean is significantly noisier than the predictions from the emulated sample. The right-hand panels of Figure A1 show the distribution of raw residuals compared to an average estimate from the emulator. [^8] The scatter about the shape parameters can vary as a function of $\bar{e}_{w}$ in both the raw dataset and emulator; here, we only compare the full distribution of residuals against the average of the standard deviations from the emulator. There is some level of disagreement in the tails of the distributions (in particular in $\gamma$), which, again, could be more effectively captured by a more sophisticated architecture, but for the most part we find that the standard deviation of the datasets match quite well with the raw simulation data.
@@
Appendix B (Lack of) Sensitivity to Weighting Scheme
Figure B1: Density Profiles with and without Weighting Scheme. Same as the left panel of Figure 2 with (left) and without (right) the weighting scheme from Rose et al. (2025b); see also description in Section 2.4). The inclusion of the weighting scheme does not significantly impact the core results of the paper.
As mentioned in Section 2.4, we employ the parameter-space weighting of Rose et al. (2025b) to down-weight parts of feedback parameter space that produce halos with properties highly inconsistent with observations. This weighting scheme is based on the stellar mass–halo mass relation (Wang et al., 2024) and down-weights the extreme parameter variations of $\bar{e}_{w}$ and $\kappa_{w}$, as well as low $\epsilon_{f,\,\mathrm{high}}$ values. Indeed, as we show in, e.g., Figure 6, both the stellar mass of the central galaxies and their supermassive black holes scale strongly with these parameters. This appendix quantifies the extent to which the resulting density profiles depend on this weighting scheme.
Figure B1 shows the density profiles with (left) and without (right) the weighting scheme applied. On the whole, there is minimal difference between the density profiles with or without the weighting. The majority of the difference comes from an increase in the central scatter (though it is only by $0.06$ dex, $20\%$, increase). This is likely due to the fact that, as we show in Figure 2, the extreme $\bar{e}_{w}$ and $\kappa_{w}$ variations drive the (relatively minor) deviations in the central density at radii $\lesssim 0.01R_{200}$. Weighting against the halos with these strong feedback parameters therefore shrinks the scatter.
Appendix C Assessing the Role of $\epsilon_{f,\,\mathrm{high}}$
@@
As mentioned in the main text, there are notable trends between physical properties of the halo and the AGN feedback. Section 3.3 lays out an argument for these trends based on an approximation of the total feedback energy (by integrating Equation 6; see also a similar argument made by Ni et al. 2023). This appendix provides supplementary support for this interpretation.
Figure C1 shows our approximation for the total feedback energy imparted by the central black hole (via Equation 17; open circles). In addition, we train a series of emulators (see Section 2.3) and report the aggregate prediction for the mean and standard deviation as the solid line and shaded region, respectively. The total feedback energy over the lifetime of the black hole is virtually unchanged $(\sim 10^{60}~{\rm erg})$ by variations in $\epsilon_{f,\,\mathrm{high}}$. It decreases slightly for feedback values less than the fiducial TNG $(\log(\epsilon_{f,\,\mathrm{high}})\lesssim-1)$, but it is completely flat above this value. This suggests that the total work that the AGN does (in the thermal mode) is roughly constant as a function of $\epsilon_{f,\,\mathrm{high}}$.
@@
Appendix D Adiabatic Contraction Fit Parameters
@@
Figure D1: Distribution of $\mathbf{A}$ and $\mathbf{w}$ parameters from DREAMS simulations. Predictions from our emulators for the best-fit AA and $w$ values (from Equation 18) as a function of $\bar{e}_{w}$ (AA and $w$ do not depend sensitively on any other DREAMS simulation parameter). The solid line shows the mean while the shaded region represents the halo-to-halo variations. The stars correspond to the average values from Hussein et al. (2025; $A=0.40$ and $w=0.45$), which fall well within the halo-to-halo variation in DREAMS.
@@
As mentioned in Section 4.2, the Gnedin et al. (2004) adiabatic contraction model includes a best-fit parameterization of the orbit-averaged radius $\bar{r}$ (Equation 18). In our work, the free parameters, $A$ and $w$ (corresponding to a normalization and exponent of the fit), are determined on a halo-by-halo basis. Not every work follows this practice; for example, Hussein et al. (2025) fit their halos using the average $A$ and $w$ parameters. We opt against this because $A$ and $w$ depend on the physics variations that DREAMS employs. In particular, they both depend on $\bar{e}_{w}$ (Figure D1). With increasing $\bar{e}_{w}$, both $A$ and $w$ decrease (though the former decreases more than the latter). These trends follow fairly closely to those observed with the central mass growth in Figure 5, i.e., halos that undergo more contraction have larger $A$ and $w$ values, which physically correspond to systematically smaller orbit-averaged radii, indicating that dark matter has moved deeper into the potential well.
@@
If we instead use the average values of $A=0.848$ and $w=0.675$, our results are qualitatively unchanged. [^9] These average values of $A$ and $w$ are different than those reported in Hussein et al. (2025; $A=0.40$ and $w=0.45$), who use six halos in the TNG model (albeit without the physics variations). However, we find that the Hussein et al. (2025) values are consistent with our results within the halo-to-halo variation (shown as the stars in Figure D1). Figure D2 shows the resulting adiabatic contraction profiles when fixing $A$ and $w$. On average, there is slightly worse agreement between the adiabatic contraction model and the true hydrodynamic simulations, but the level to which the two disagree is within $\sim 10\%$. The (weighted) scatter also increases when considering fixed $A$ and $w$. Our key conclusion regarding the comparison with the bursty FIRE feedback model remains qualitatively unchanged, however. Even when ignoring the variations of $A$ and $w$ due to the physics variations, we find that the deviations seen in FIRE halos are significantly larger than those of the DREAMS halos.
Figure D2: Ratio of Mass Contraction to Hydrodynamic Simulation Mass Profiles using fixed $\mathbf{A}$ and $\mathbf{w}$. Same as Figure 9, but assuming a fixed $A$ and $w$ value that corresponds to the average of the DREAMS distribution ($A=0.848$ and $w=0.675$). We still find qualitatively different behavior between FIRE-2 and DREAMS.

### v2

@@ IV.2.1 Comparison with Bursty Feedback Model
It is not clear, however, whether bursty feedback disrupting the potential and preventing adiabatic contraction is a generic feature of FIRE-2 or a feature of the specific implementation of the model. An analogous suite of Milky Way-mass halos within the FIRE-2 model (or its successor FIRE-3; Hopkins et al. 2023) would be particularly valuable for constraining this. Such a comparison would help determine whether the deviations from simple adiabatic contraction models identified by Hussein et al. (2025) are a generic outcome of bursty feedback or depend sensitively on the details of how feedback is modeled in FIRE.
V Conclusion
@@
  • Overall, there are only minor variations in the density profiles of Milky Way-mass halos with variations to the IllustrisTNG model (Section III.1 Figure 2). Virtually all variation in the density profiles is driven by halo-to-halo variation, with a relatively small contribution from the two supernova parameters that only create differences of $\pm\sim 20-50\%$ in the inner regions ($\lesssim 0.01R_{200}$) of the halos.
@@
  • We fit the density profiles with a generalized NFW profile (e.g., Jaffe, 1983; Hernquist, 1990) which has two normalization parameters and three shape parameters (Equation 9, Section III.2). Again, intrinsic halo-to-halo variation plays the dominant role in setting both the normalization and shape of halos. Beyond the halo-to-halo variance, changes to the supernova wind speed ($\kappa_{w}$) can also vary the normalization of the profile (Figure 3) and increases to the wind energies ($\bar{e}_{w}$) can reduce the deviations from an NFW profile (Figure 4). Physically, this suggests that ejecting faster winds more strongly disrupts the potential, whereas more wind energy effectively suppresses stellar mass growth.
@@
  • We quantify the impact of baryons on the dark matter halos by defining the central mass growth as ratio of the DMO simulation enclosed mass to the hydro simulation enclosed mass at $0.01R_{200}$ (Figure 5). Here, the astrophysics variations play a dominant role. Larger values of $\bar{e}_{w}$ and $\kappa_{w}$ are more efficient at preventing stellar mass growth (left two panels of Figure 6), thus the halo contracts less. On the other hand, we find the opposite trend in $\epsilon_{f,\,\mathrm{high}}$: increasing black hole feedback increases the work required to contract (right panel of Figure 6, see also Appendix B).
@@
  • We compared the results of the hydro simulations to DMO simulations with matched initial conditions (Section IV.1). On the whole, the DMO simulations have a lower density in the inner regions ($\lesssim 0.01R_{200}$) than the hydro simulations (Figure 7). The extent to which the density is enhanced depends on the feedback, however. In fact, higher values of $\bar{e}_{w}$ tend towards a ratio of unity with respect to the DMO simulations, in good agreement with the results of the profile shape and central mass growth.
@@
  • We quantified whether or not our halos are consistent with approximations to adiabatic contraction (Section IV.2). Halos generally agree within $\sim 10\%$ regardless of the feedback or cosmology (Figure 8); however, systems with late-time mergers that occur at slightly different times between the hydro and DMO simulations tend to disagree significantly. As with many previous results, the supernova wind energy changes the picture due to the lack of baryon mass accumulation: these halos contract adiabatically.Finally, we showed that the bursty feedback FIRE(-2) model produces significantly larger deviations from adiabatic contraction than the average DREAMS halos (Figure 9). This suggests that what drives deviations from adiabatic contractions are bursts of stellar feedback, not just strong feedback.
@@
Taken together, these results show that halo-to-halo variance is the dominant source of scatter in Milky Way–mass dark matter profiles in the TNG model (within the DREAMS parameter variations), with $\bar{e}_{w}$ and $\kappa_{w}$ feedback parameters producing secondary but non-negligible effects. AGN feedback plays a smaller role, and cosmological parameter variations are essentially irrelevant at this mass scale.
@@
Raw data from the DREAMS CDM suite is available via Globus transfer at dreams-project.readthedocs.io/en/latest/data_access.html. The scripts used to generate the reduced data and all other scripts used to support the findings of this work are made publicly available 10.5281/zenodo.18715476 (catalog doi: 10.5281/zenodo.18715476) and on github github.com/AlexGarcia623/DREAMS-Density-Profiles
@@ Acknowledgements
AMG, PT, JK, NA, AB, XO, AF, and NK acknowledge support from the National Science Foundation under Cooperative Agreement 2421782 and the Simons Foundation grant MPS-AI-00010515 awarded to NSF-Simons AI Institute for Cosmic Origins – CosmicAI, https://www.cosmicai.org/. ML is supported by the Department of Energy (DOE) under Award Number DE-SC0007968, as well as the Simons Investigator Award. AC acknowledges the hospitality of the Weizmann Institute of Science and the support from the Benoziyo Endowment Fund for the Advancement of Science. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation programme (grant agreement No. 101117510). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union. The European Union cannot be held responsible for them. LN is supported by the Sloan Fellowship, the NSF CAREER award 2337864, NSF award 2307788, and by the NSF award PHY2019786 (The NSF AI Institute for Artificial Intelligence and Fundamental Interactions, http://iaifi.org/) XS acknowledges the support of the NASA theory grant JWST-AR-04814. This work was performed in part at Aspen Center for Physics, which is supported by National Science Foundation grant PHY-2210452. The authors also thank the Simons Foundation for their support in hosting and organizing workshops on the DREAMS Project. HL is supported by the U.S. Department of Energy under grant DE-SC0026297 and the Cecile K. Dalton Career Development Professorship, endowed by Boston University trustee Nathaniel Dalton and Amy Gottleib Dalton.
@@ Appendix A Emulator Validation Test
This work employs a series of neural network emulators to coherently interpret the DREAMS data. We validate each emulator (Figures 3456C1, and D1) by comparing its predictions to the raw simulation outputs (following R25a). Figure A1 shows an example of this validation process. The left panel of Figure A1 is identical to Figure 4 with the addition of the true rolling mean (solid black line) and standard deviation (gray shaded region) from the measured simulation data points. The raw simulation results largely track the emulator ones, although there is a small shift in the outer slope. One potential reason for this is that the emulator is trained to predict single parameter dependencies (i.e., we keep all other parameters fixed), whereas the raw simulation outputs include inherent co-variation of parameters. To capture this co-variance, we would need a different architecture, such as normalizing flows (see, e.g., Lilie et al., 2025). Regardless, the emulated prediction for the mean is largely consistent with these raw simulation outputs even with our simple emulator. Figure A1 further demonstrates the utility of the emulator: the true rolling mean is significantly noisier than the predictions from the emulated sample. The right-hand panels of Figure A1 show the distribution of raw residuals compared to an average estimate from the emulator. [^9] The scatter about the shape parameters can vary as a function of $\bar{e}_{w}$ in both the raw dataset and emulator; here, we only compare the full distribution of residuals against the average of the standard deviations from the emulator. There is some level of disagreement in the tails of the distributions (in particular in $\gamma$), which, again, could be more effectively captured by a more sophisticated architecture, but for the most part we find that the standard deviation of the datasets match quite well with the raw simulation data.
@@
Appendix B Assessing the Role of $\epsilon_{f,\,\mathrm{high}}$
@@
There are notable trends between physical properties of the halo and the AGN feedback. Section III.3 lays out an argument for these trends based on an approximation of the total feedback energy (by integrating Equation 6; see also a similar argument made by Ni et al. 2023). This appendix provides supplementary support for this interpretation.
Figure C1 shows our approximation for the total feedback energy imparted by the central black hole (via Equation 17; open circles). In addition, we train a series of emulators (see Section II.3) and report the aggregate prediction for the mean and standard deviation as the solid line and shaded region, respectively. The total feedback energy over the lifetime of the black hole is virtually unchanged $(\sim 10^{60}~{\rm erg})$ by variations in $\epsilon_{f,\,\mathrm{high}}$. It decreases slightly for feedback values less than the fiducial TNG $(\log(\epsilon_{f,\,\mathrm{high}})\lesssim-1)$, but it is completely flat above this value. This suggests that the total work that the AGN does (in the thermal mode) is roughly constant as a function of $\epsilon_{f,\,\mathrm{high}}$.
@@
Appendix C Adiabatic Contraction Fit Parameters
@@
Figure D1: Dependence of $\mathbf{A}$ and $\mathbf{w}$ on DREAMS feedback variations. Left: Predictions from our emulators for the best-fit $A$ and $w$ values (from Equation 18) as a function of $\bar{e}_{w}$ (AA and $w$ do not depend sensitively on any other DREAMS simulation parameter). The solid line shows the mean while the shaded region represents the halo-to-halo variations. The stars correspond to the average values from Hussein et al. (2025; $A=0.40$ and $w=0.45$), which fall well within the halo-to-halo variation in DREAMS. Right: Same as Figure 9, but assuming a fixed $A$ and $w$ value that corresponds to the average of the DREAMS distribution ($A=0.848$ and $w=0.675$). We still find qualitatively different behavior between FIRE-2 and DREAMS.
@@
As mentioned in Section IV.2, the Gnedin et al. (2004) adiabatic contraction model includes a best-fit parameterization of the orbit-averaged radius $\bar{r}$ (Equation 18). In our work, the free parameters, $A$ and $w$ (corresponding to a normalization and exponent of the fit), are determined on a halo-by-halo basis. Not every work follows this practice; for example, Hussein et al. (2025) fit their halos using the average $A$ and $w$ parameters. We opt against this because $A$ and $w$ depend on the physics variations that DREAMS employs. In particular, they both depend on $\bar{e}_{w}$ (left panel of Figure D1). With increasing $\bar{e}_{w}$, both $A$ and $w$ decrease (though the former decreases more than the latter). These trends follow fairly closely to those observed with the central mass growth in Figure 5, i.e., halos that undergo more contraction have larger $A$ and $w$ values, which physically correspond to systematically smaller orbit-averaged radii, indicating that dark matter has moved deeper into the potential well.
@@
If we instead use the average values of $A=0.848$ and $w=0.675$, our results are qualitatively unchanged. [^10] These average values of $A$ and $w$ are different than those reported in Hussein et al. (2025; $A=0.40$ and $w=0.45$), who use six halos in the TNG model (albeit without the physics variations). However, we find that the Hussein et al. (2025) values are consistent with our results within the halo-to-halo variation (shown as the stars in the left panel of Figure D1). The right-hand panel of Figure D1 shows the resulting adiabatic contraction profiles when fixing $A$ and $w$. On average, there is slightly worse agreement between the adiabatic contraction model and the true hydrodynamic simulations, but the level to which the two disagree is within $\sim 10\%$. The (weighted) scatter also increases when considering fixed $A$ and $w$. Our key conclusion regarding the comparison with the bursty FIRE feedback model remains qualitatively unchanged, however. Even when ignoring the variations of $A$ and $w$ due to the physics variations, we find that the deviations seen in FIRE halos are significantly larger than those of the DREAMS halos.